OpenStudy (anonymous):

How do you factor (p+q)^3-(p-q)^3? Please explain, don't just give the answer.

OpenStudy (jamesj):

First, do you know how to factor in general the expression: a^3 - b^3 ?

OpenStudy (anonymous):

(a-b) (a^2+a b+b^2)

OpenStudy (jamesj):

Correct. Now set a = p + q and b = .... and apply your formula.

OpenStudy (anonymous):

Does b equal p-q? It got cut off

OpenStudy (jamesj):

I didn't say, on purpose. But that would seem to make sense, wouldn't it?

OpenStudy (anonymous):

Ah I see.

OpenStudy (jamesj):

Happy factoring!

OpenStudy (jamesj):

Just make sure you check your signs as you go.

OpenStudy (anonymous):

I got 2 (3 p^2 q+q^3)

OpenStudy (jamesj):

Looks good to me, as there are lots of cancelations.

OpenStudy (anonymous):

Thanks again.

OpenStudy (jamesj):

if we just calculated it out (p+q)^3 - (p-q)^3 = p^3 + 3p^2q + 3pq^2 + q^3 -p^3 + 3p^2q - 3pq^2 + q^3 = 0 + 6p^2q + 0 + 2q^3 = 2(3p^2q + q^3)

OpenStudy (jamesj):

So, likes lots of this in mathematics, there's more than one way to skin a cat. Anyway -- got to move on! Cheers.